OpenStudy (anonymous):

Write the slope-intercept equation for the line with the given slope that contains the given point. m= -3/4; (0,5)

OpenStudy (anonymous):

y = (-3/4)x + 5

OpenStudy (anonymous):

now medal me you wingspanlogging moron

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

Point-Slope Formula: y-y1=m(x-x1) where m is the slope, and (x1,y1) is the given point lets use the Point-Slope Formula to find the equation of the line y-5=(-3/4)(x-0) Plug in m=-3/4, x1=0 , y1=5 and (these values are given) and so on ...